# Thickness-dependent conductivity of nanometric semiconductor thin films

### v1

@@ Abstract
Figure 1: Panel (a) shows the thin film geometry in real space (confined along the $z$-axis but unconfined along the $x$ and $y$ axis), with the maximum wavelength of a free carrier that corresponds to a certain polar angle $\theta$. Panel (b) shows the corresponding geometry of electronic states in $k$-space, where the outer Fermi sphere (of radius $k_{F}$) contains two symmetric spheres of hole pockets (states suppressed by confinement), i.e. states in $k$-space that remain unoccupied due to confinement along the $z$-axis. See Refs. [9, 10, 11] for a detailed mathematical derivation of this result.
As derived in Ref. [10], plane-wave electronic states of free carriers with wavelength $\lambda$ larger than $\lambda_{max}=\frac{L}{\cos\theta}$ cannot propagate in the thin film along the direction $\theta$ (cfr. Fig. 1(a)), where the polar angle $\theta$ of the propagation direction is measured with respect to the $z$ axis (cfr. Fig. 1(a)). The momenta $k_{x}$, $k_{y}$, $k_{z}$ are, in general, discretized for small systems if one, arbitrarily, chooses vanishing (hard-wall) boundary conditions (BCs) at the interfaces for the wavefunctions of the governing Schrödinger equation. However, if the sample is extended in the $xy$ plane (for at least several microns in both directions), as it is for thin films, $k$ can still be treated as a quasi-continuous variable [10, 12], because $|\mathbf{k}|=k=2\pi/\lambda$ obeys the following relation [13, 14]:
@@
The confinement-induced cutoff on $\lambda$ means that a number of electronic states in $k$-space are suppressed due to the confinement along the $z$ direction of the film. For a film of thickness $L$, this reduction of the available volume for free carriers in momentum ($k$) space is evaluated exactly as (cfr. Fig. 1(b)):
@@
We consider a c-Si semiconductor thin film, which is either intrinsic or weakly-doped such as e.g. the ex situ-doped thin films studied recently in Ref. [8]. For these systems, the concentration of free carriers cannot be precisely determined, and will vary broadly from a lower bound that coincides with the intrinsic material, $n\sim 10^{16}$ m-3, to an upper bound of $n\sim 10^{25}$ m-3 [8]. Here we focus on a regime of weak n-doping where $n\sim 10^{16}-10^{20}$m-3. In these conditions, $L_{c}$ is very large (in the order of microns). Since we are interested in nanometric thin films ($L=1-100$ nm), we can thus safely operate in the regime $L<L_{c}$, using the second of the two relations in Eq. (4).
@@
A standard contribution to the resistivity in thin films and other confined materials is due to electron scattering with the interface, or surface scattering. This contribution was computed by Fuchs [16] and Sondheimer [17] using the Boltzmann equation framework. As reported by Sondheimer, simple closed-form expressions for the resistivity contribution of surface scattering are obtained for thick films and thin films, respectively, as $\rho_{s}/\rho_{0}=[1+3/(8\kappa)]^{-1}$ and $\rho_{s}/\rho_{0}=\{4/[3\kappa\ln(1/\kappa)]\}^{-1}$, with $\kappa=L/\ell$ where $\ell$ is the mean free path in the bulk material [17]. For silicon, $\ell\approx 20$ nm. The crossover between the two formulae occurs, therefore, around $10$ nm. Assuming, as usual, that Matthiessen’s rule applies, the different effects of quantum confinement and of surface scattering can be added as independent contributions [13]. We then have the two thickness-dependent contributions, from quantum wave confinement and from the Fuchs-Sondheimer (FS) scattering, respectively, summing up (Matthiessen’s rule) to give the total resistivity $\rho(L)$ as:
@@
Figure 2: Comparison between the theoretical predictions of the proposed model (solid continuous line) obtained by combining the FS surface scattering theory with the electronic confinement model via Matthiessen’s rule Eq. (11). The green dashed line represents the FS surface scattering prediction without the electron-confinement correction, whereas the orange dashed line represents the electron confinement correction. The symbols (circles) are the experimental data from Ref. [8].
@@
In summary, we have developed a microscopic theory of confinement-induced conductivity and resistivity in semiconductor thin films which is able to describe and explain the experimental data on c-Si nanometric thin films down to a thickness of about $4.5$ nm. Without the theory presented in this paper, it would be impossible to explain the sharp increase in resistivity, by one order of magnitude, observed experimentally [8] when the film thickness goes below $10$ nm. The theoretical analysis shows that there is a fundamental crossover, at about $10$ nm for the case of weakly-doped c-Si, from a regime (at larger thickness) dominated entirely by the Fuchs-Sondheimer surface-scattering contribution to the resistivity, to a new regime, below $10$ nm, where the resistivity is, instead, entirely dominated by quantum confinement. These results will pave the way for a rational understanding and design of new generation microchips with nanometer-scale thickness, within the frame of the ”2nm process” [5], and will find broad applications from quantum gated logic and quibts design to solar energy.

### v2

@@ Abstract
Figure 1: Panel (a) shows the thin film geometry in real space (confined along the $z$-axis but unconfined along the $x$ and $y$ axis), with the maximum wavelength of a free carrier that corresponds to a certain polar angle $\theta$. Panel (b) shows the corresponding geometry of electronic states in $k$-space, where the outer Fermi sphere (of radius $k_{F}$) contains two symmetric spheres of hole pockets (states suppressed by confinement), i.e. states in $k$-space that remain unoccupied due to confinement along the $z$-axis. In (c), for strong confinement (e.g. quasi-2D films), the hole pockets of forbidden states have grown to the point that the Fermi surface gets significantly distorted into a surface belonging to a different homotopy group $\mathbb{Z}$. See Refs. [9, 10, 11] for a detailed mathematical derivation of these results. See Refs. [9, 10, 11] for a detailed mathematical derivation of this result.
As derived in Ref. [10], plane-wave electronic states of free carriers with wavelength $\lambda$ larger than $\lambda_{max}=\frac{L}{\cos\theta}$ cannot propagate in the thin film along the direction $\theta$ (cf. Fig. 1(a)), where the polar angle $\theta$ of the propagation direction is measured with respect to the $z$ axis (cf. Fig. 1(a)). The momenta $k_{x}$, $k_{y}$, $k_{z}$ are, in general, discretized for small systems if one, arbitrarily, chooses vanishing (hard-wall) boundary conditions (BCs) at the interfaces for the wavefunctions of the governing Schrödinger equation. However, if the sample is extended in the $xy$ plane (for at least several microns in both directions), as it is for thin films, $k$ can still be treated as a quasi-continuous variable [10, 12], because $|\mathbf{k}|=k=2\pi/\lambda$ obeys the following relation [13, 14]:
@@
The confinement-induced cutoff on $\lambda$ means that a number of electronic states in $k$-space are suppressed due to the confinement along the $z$ direction of the film. For a film of thickness $L$, this reduction of the available volume for free carriers in momentum ($k$) space is evaluated exactly as (cf. Fig. 1(b)):
@@
We consider a c-Si semiconductor thin film, which is either intrinsic or weakly-doped such as e.g. the ex situ-doped thin films studied recently in Ref. [8]. Hence, we are going to use equations for the free-carriers concentration as for intrinsic semiconductors. These equations can still be used even for doped semiconductors as long as the concentrations of electrons and holes are comparable [6]. Hence, it is assumed that the theoretical model developed below is applicable to thin films doped ex situ, which are well known to be affected by poor dopant incorporation and by dopant-deactivation issues, as discussed in Ref. [8], see also Ref. [16]. For nanowires, this effect was explained by simulations in terms of suppressed ionization near the surface due to electronic confinement, with the concomitant formation of image charges outside the wire [17] (cfr. [18] for applications of these effects to tailor the conductivity in nanowires). The dopant deactivation effect in ex-situ doped thin films is, however, the result of complex transport phenomena and confinement effects, which cannot be quantitatively modelled in the absence of specific measurements, that are currently out of reach for available experimental techniques. Also due to this spatially non-uniform dopant-deactivation effect, the doping in these systems is known to be spatially-dependent, with a concentration profile which is generally unknown. Since, anyway, we are interested in the overall resistivity of the sample, which is an averaged quantity over the whole film, we shall emphasize that the carrier concentration in our equations is meant to result from a spatial averaging process over an underlying, as-yet-unknown, spatial profile. While the current model does not consider that all dopants are ionized (on the contrary, given the considerations stated above), and uses an effective, spatially-averaged carrier density, this issue should be kept in mind when considering the comparison with experimental data, in what follows.
For these systems [8], the concentration of free carriers cannot be determined precisely, and will vary broadly from a lower bound that coincides with the intrinsic material, $n\sim 10^{16}$ m-3, to an upper bound of $n\sim 10^{25}$ m-3 as noted in [8]. Here we focus our theoretical analysis on a regime of very weak n-doping where $n\sim 10^{16}-10^{20}$m-3.
In these conditions, $L_{c}$ is quite large (of the order of hundreds of nanometers). Since we are interested in nanometric thin films ($L=1-100$ nm), we can thus safely operate in the regime $L<L_{c}$, using the second of the two relations reported in Eq. (4). As it shall become clearer below, we should anticipate that this $L$-dependence due to electron confinement is effectively obscured by the surface-scattering Fuchs-Sondheimer (FS) $L$-dependence down to a certain film thickness at which one observes the crossover from the FS $L$-dependence to the one predicted by electron confinement.
@@
A standard contribution to the resistivity in thin films and other confined materials is due to electron scattering with the interface, or surface scattering [19]. This contribution was computed by Fuchs [20] and Sondheimer [21] using the Boltzmann equation framework. As reported by Sondheimer, simple closed-form expressions for the resistivity contribution of surface scattering are obtained for thick films and thin films, respectively, as $\rho_{s}/\rho_{0}=[1+3/(8\kappa)]^{-1}$ and $\rho_{s}/\rho_{0}=\{4/[3\kappa\ln(1/\kappa)]\}^{-1}$, with $\kappa=L/\ell$ where $\ell$ is the mean free path in the bulk material [21]. For silicon, $\ell\approx 20$ nm. The crossover between the two formulae occurs, therefore, around $10$ nm. Assuming, as usual, that Matthiessen’s rule applies, the different effects of quantum confinement and of surface scattering can be added as independent contributions [13]. We then have the two thickness-dependent contributions, from quantum wave confinement and from the Fuchs-Sondheimer (FS) scattering, respectively, summing up (Matthiessen’s rule) to give the total resistivity $\rho(L)$ as:
@@
Figure 2: Comparison between the theoretical predictions of the proposed model (solid continuous line) obtained by combining the FS surface scattering theory with the electronic confinement model (Eq. (10)) via Matthiessen’s rule, Eq. (11). The green dashed line represents the FS surface scattering prediction (see text for the corresponding equations) without the electron-confinement correction, whereas the orange dashed line represents the electron confinement correction without the FS contribution, given by Eq. (10). The symbols (circles) are the experimental data from Ref. [8].
@@
Regarding the application of the present theory to the experimental data, the following considerations are in order: (i) the theoretical model applies most naturally to the top data points in Fig. 2, which are indeed well fitted by the solid curve (theoretical prediction), since they refer to poorly-doped samples (weakly doped or nearly-intrinsic regime) [8]. These data points exhibit the upturn in resistivity below 10 nm because, indeed, the physics is controlled by electronic confinement, as explained by the model, Eq. (10), which predicts an exponential increase of resistivity with decreasing $L$ in that range. (ii) the data points that lie at the bottom (i.e. with, comparatively, the lowest resistivity values), as explained in Ref. [8], correspond to higher active-donors concentrations. Hence, for these data, the mechanism is extrinsic. This consideration is supported by the fact that their resistivity is almost insensitive to the thickness $L$. The proposed theoretical model does not apply to these low-lying data points, but, for completeness, it is important to show them as well as they were reported in Ref. [8].
Regarding the more general applicability of the proposed model, a few considerations are necessary. The model will be most useful in order to optimize the resistivity, and, hence, the energy and signal transmission efficiency, in materials where the free carrier concentration is low, such that the second of Eq. (4) applies. This is the case of semiconductors (Si, Ge, etc) as well as semi-metals, including quasi-2D semiconductors like molybdenum disulfide and topological semi-metals [19]. In these systems, the equations presented above can be used to tune the resistivity of sub-10 nm films by varying the carrier concentration, e.g. by controlling the doping. For good metals, like Cu, we are, instead, in a regime where the first of Eq. (4) applies, and therefore this effect is smaller and probably negligible with respect to the FS surface-scattering effect [22]. Hence, Eq. (4) together with the model presented here, can be used by practitioners to tune and optimize the energetic and signal transmission performance of a given sub-10 nm material of the categories mentioned above, for a certain device.
In summary, we have developed a microscopic theory of confinement-induced conductivity and resistivity in semiconductor thin films which is able to describe and explain the experimental data on c-Si nanometric thin films down to a thickness of about $4.5$ nm. Without the theory presented in this paper, it would be impossible to explain the sharp increase in resistivity, by one order of magnitude, observed experimentally [8] when the film thickness goes below $10$ nm. The theoretical analysis shows that there is a fundamental crossover, at about $10$ nm for the case of weakly-doped c-Si, from a regime (at larger thickness) dominated entirely by the Fuchs-Sondheimer surface-scattering contribution to the resistivity, to a new regime, below $10$ nm, where the resistivity is, instead, entirely dominated by quantum confinement. Interestingly, the same upturn in resistivity with decreasing film thickness below about $10$ nm has been recently observed in first-principles simulations of a very different system such ultrathin copper films [22], which points to a certain degree of universality in the predicted increase of resistivity due to quantum confinement in ultrathin films. These results will pave the way for a rational understanding and design of new generation microchips with nanometer-scale thickness, within the frame of the ”2nm process” [5], and will find broad applications from quantum gated logic and qubits design to solar energy.